# Rethinking the Role of Prompting Strategies in LLM Test-Time Scaling: A Perspective of Probability Theory

### v1

@@ Abstract
Recently, scaling test-time compute on Large Language Models (LLM) has garnered wide attention. However, there has been limited investigation of how various reasoning prompting strategies perform as scaling. In this paper, we focus on a standard and realistic scaling setting: majority voting. We systematically conduct experiments on 6 LLMs $\times$ 8 prompting strategies $\times$ 6 benchmarks. Experiment results consistently show that as the sampling time and computational overhead increase, complicated prompting strategies with superior initial performance gradually fall behind simple Chain-of-Thought. We analyze this phenomenon and provide theoretical proofs. Additionally, we propose a method according to probability theory to quickly and accurately predict the scaling performance and select the best strategy under large sampling times without extra resource-intensive inference in practice. It can serve as the test-time scaling law for majority voting. Furthermore, we introduce two ways derived from our theoretical analysis to significantly improve the scaling performance. We hope that our research can promote to re-examine the role of complicated prompting, unleash the potential of simple prompting strategies, and provide new insights for enhancing test-time scaling performance.
@@ 2.1 Models
We conduct experiments on 4 open-sourced LLMs including Qwen2.5-7B-Instruct (Yang et al., 2024a), LLaMA-3-8B-Instruct (Dubey et al., 2024), GLM-4-9B-Chat (GLM et al., 2024) and Phi-3.5-mini-Instruct, and 2 close-sourced LLMs including Gemini-1.5-Flash (Team et al., 2024) and GPT-4o-mini (Open AI, 2024a).
@@ 4 Why CoT Performs worse with Lower $N$ while better with Larger $N$?
$\begin{split}\mathcal{X}_{ir}=\{\mathbf{x}_{i,j}\,|\,\mathbf{x}_{i,j}=\max\{\mathbf{X}_{i}\}\}\\ k\sim\text{Uniform}(\mathcal{X}_{ir}),\quad{a}^{*}_{i}={a}_{k}~{}~{}~{}\end{split}$ (4)
@@
Table 1: Difficulty proportion of questions and extreme peformance (denote by “Acc”) for each $\mathbf{P}_{i}$ and LLM across benchmarks. CoT has more easy questions and fewer hard questions, and can reach the best extreme performance on all LLMs. More results on other models are shown in Table 4.
@@
Table 2: Quantity of questions described in Section 4.2. Results are displayed as “x/y”, where x is for Qwen and y for LLaMA. The value $v_{ij}$ in the $i$th row and $j$th column represents the quantity of data that satisfies Theorem 4. Results prove that CoT has greater potential to significantly increase performance as scaling.
@@
DiP $\uparrow$ -
447/816
414/794
457/459
393/513
1711/2582
CoT $\uparrow$
423/620
374/646
416/382
361/408
1574/2046
L2M $\uparrow$
505/639
510/677
494/432
403/393
1912/2141
SBP $\uparrow$
599/1316
601/1433
564/1398
429/923
2193/5070
AnP $\uparrow$
800/1243
817/1381
799/1380
776/871
3192/4875
Sum $\uparrow$
2327/3818
2375/4307
2151/4218
2143/2144
1586/2237
@@ 4.1 CoT Has More Easy Questions and Fewer Hard Questions
We identify two primary reasons why CoT sometimes performs worse with lower sample sizes ($N$) but achieves better performance among these prompting approaches with larger $N$. The first reason relates to the distribution of question difficulty for CoT. CoT has more easy questions and fewer hard questions. When sampling with lower $N$, $\mathbf{P}_{i}$ still has a small probability of obtaining the right answer for hard questions, while the probability diminishes to zero as increasing $N$. This is the opposite of easy questions. The prompting strategy with fewer hard questions and more easy questions will improve performance more rapidly when scaling. According to Theorem 1 to 3, we can calculate the extreme performance of $\mathbf{P}_{i}$ according to the difficulty proportion of questions, i.e., $\sum\limits_{x\in\mathfrak{D}}\lim\limits_{N\to+\infty}\!\!\!\!\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$. Table 1 summarizes the difficulty proportion of the questions and extreme performance for each $\mathbf{P}_{i}$ on each model. It can be observed that CoT has more easy questions and fewer hard questions, and can reach the best extreme performance on all models, thus making CoT gradually dominate as increasing $N$ even if it has a lower pass@1 accuracy.
@@ 5 Predicting Scaling Performance and $\mathbf{P}^{*}_{N}$
In practice, evaluating the test-time scaling performance requires significantly intensive resource consumption, especially with very large sampling time $N$. For pretraining, it is feasible to predict the train-time scaling performance based on the scaling law (Kaplan et al., 2020) through a series of low-cost experiments, while maintaining the model architecture largely unchanged and minimizing the risks associated with large-scale training. Similarly, we can also use the sample results of $\mathbf{P}_{i}$ with fewer $N$ to approximately get the distribution $\{p_{i,1},p_{i,2},\dots,p_{i,m}\}$ to predict the test-time scaling performance with larger $N$. Directly, one can utilize the multinomial distribution probability calculation formula (Equation 13 and 14) to calculate $\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ with enumeration or leverage numerical simulation to estimate. However, their computational complexities are both $O(N)$, and the former needs to traverse all situations and is difficult to operate. Therefore, we propose a method with the computational complexity $O(1)$ to quickly predict the scaling performance of majority voting for arbitrary $\mathbf{P}_{i}$, which can serve as the test-time scaling law. It can also select the best prompting strategy $\mathbf{P}^{*}_{N}$ according to the predicted performances of each $\mathbf{P}_{i}$.
Here we omit the prompting index $i$ and input question $x$, and assume ${a}_{1}$ is the correct answer in the following. According to Khinchin’s Law of Large Numbers and Lindeberg-Levy Central Limit Theorem, when $N$ is large enough, each $\mathbf{x}_{j}$ can be approximated by a normal distribution. Specifically, for $\mathbf{x}_{1}$, we have
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!E[Z]=N(p_{1}-p_{max}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$Var[Z]=N(p_{1}(1-p_{1})+p_{max}(1-p_{max})),$
$\mathbf{P}^{*}_{N}$ Sampling Time $N$
1 3 5 10 20 50 100 1000
Real L2M CoT CoT CoT SBP SBP SBP SBP
Predicted L2M CoT CoT CoT SBP SBP SBP SBP
Correct
@@ 7 Related Work
CoT series carefully design exemplars or 0-shot prompts to unleash the potential of step-by-step solving (Wei et al., 2022; Kojima et al., 2022; Zhang et al., 2023; Fu et al., 2023). Zhou et al. (2023); Dua et al. (2022); Khot et al. (2023) break down the question into smaller, more manageable subproblems. Madaan et al. (2023); Kim et al. (2023) force LLMs to self-evaluate and correct. Du et al. (2024); Liang et al. (2024); Chan et al. (2024); Liu et al. (2025) utilize multi-agent debate to collaborate reasoning. Yasunaga et al. (2024); Yu et al. (2024) guide LLMs to draw experience from analogous problems. Zheng et al. (2024); Gao et al. (2025) promote LLMs on abstract reasoning.
@@ 8 Conclusion
We comprehensively study the behavior of various prompting strategies when scaling majority voting. Our experiments on 6 LLMs $\times$ 8 prompting strategies $\times$ 6 benchmarks consistently show that CoT has the potential to perform best as scaling. Theoretical analyses reveal that CoT benefits from fewer hard questions, more easy questions, and less susceptibility to incorrect answers, enabling rapid performance gains with increased sampling. Additionally, our proposed method for predicting scaling performance offers a practical tool to select the optimal prompting strategy under given budgets. What’s more, we introduce two effective methods to further improve scaling performance. We hope our research can provide new insights to understand, predict, and improve LLM test-time scaling performance.
@@ Appendix A Broader Related Work
Several studies have shown that scaling test-time compute can be more effective than scaling model parameters (Snell et al., 2025; Open AI, 2024b). Aggarwal et al. (2023); Li et al. (2024); Chen et al. (2024a) improve the reasoning efficiency with majority voting by adjusting the sampling time. Damani et al. (2025); Zhang et al. (2024b) learn to dynamically allocate resources under limited sampling time budgets. Chen et al. (2024b); Yue et al. (2024); Šakota et al. (2024) leverage multiple models with different prices to reduce cost while maintaining performance.
@@
Jin et al. (2024b) studies the impact of reasoning step length of CoT. Wang et al. (2023a) studies what makes CoT prompting effective, indicating that being relevant to the query and correctly ordering the reasoning steps are more important. Feng et al. (2024); Cui et al. (2024) analyze the mechanism of CoT from a theoretical perspective. Sprague et al. (2025) points out that CoT helps mainly on math and symbolic reasoning by sorting and analyzing a large number of experimental results. Chen et al. (2024c) proposes a framework to quantify the reasoning boundary of CoT. Yang et al. (2024b) provides an in-context learning analysis of CoT. Chen et al. (2024a) investigates and analyzes the performance changes with more LLM calls. Chen et al. (2024d) proves that the failure probability of test-time scaling decays to zero exponentially or by a power law.
@@ Appendix B Proofs
$\!\!\!\!\!\!p_{1}^{k}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{t-1}^{k_{t-1}}p_{t}^{k_{t}+1}p_{t+1}^{k_{t+1}}\cdots p_{m}^{k_{m}}$. It can be seen that case 1 and case 3 are also present in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$, whereas case 2 does not. We begin by considering case 1 and case 2. The terms in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ corresponding to case 1 $p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}$ are shown in Equation 19, and no term in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ involves case 2. The terms in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ involving case 1 are shown in Equation 20. The terms in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ involving case 2 $p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k+1}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}$ are shown in Equation 21. Based on the fact of Equation 22, we can establish the inequality Equation 23, i.e., the terms corresponding to case 1 and case 2 in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ are greater than or equal to those in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$.

### v2

@@ Abstract
Recently, scaling test-time compute on Large Language Models (LLM) has garnered wide attention. However, there has been limited investigation of how various reasoning prompting strategies perform as scaling. In this paper, we focus on a standard and realistic scaling setting: majority voting. We systematically conduct experiments on 6 LLMs $\times$ 8 prompting strategies $\times$ 6 benchmarks. Experiment results consistently show that as the sampling time and computational overhead increase, complicated prompting strategies with superior initial performance gradually fall behind simple Chain-of-Thought. We analyze this phenomenon and provide theoretical proofs. Additionally, we propose a probabilistic method to efficiently predict scaling performance and identify the best prompting strategy under large sampling times, eliminating the need for resource-intensive inference processes in practical applications. Furthermore, we introduce two ways derived from our theoretical analysis to significantly improve the scaling performance. We hope that our research can promote to re-examine the role of complicated prompting, unleash the potential of simple prompting strategies, and provide new insights for enhancing test-time scaling performance. Code is available at https://github.com/MraDonkey/rethinking_prompting.
@@ 2.1 Models
We conduct experiments on 4 open-sourced LLMs including Qwen2.5-7B-Instruct (Yang et al., 2024a), LLaMA-3-8B-Instruct (Dubey et al., 2024), GLM-4-9B-Chat (GLM et al., 2024) and Phi-3.5-mini-Instruct, and 2 closed-sourced LLMs including Gemini-1.5-Flash (Team et al., 2024) and GPT-4o-mini (Open AI, 2024a).
@@ 4 Why CoT Performs worse with Lower $N$ while better with Larger $N$?
$\begin{split}\mathcal{J}_{i}=\{j\,|\,\mathbf{x}_{i,j}=\max\{\mathbf{X}_{i}\}\}\\ k\sim\text{Uniform}(\mathcal{J}_{i}),\quad{a}^{*}_{i}={a}_{k}\end{split}$ (4)
@@
Phi-3.5-mini-Instruct Gemini-1.5-Flash GPT-4o-mini
DiP 78.4% 0.6% 21.1% 78.6 91.0% 0.0% 9.0% 91.0 89.7% 0.4% 9.9% 89.9
CoT 81.2% 0.4% 18.4% 81.4 91.2% 0.2% 8.6% 91.3 89.8% 0.3% 9.9% 90.0
L2M 80.2% 0.6% 19.2% 80.5 90.9% 0.2% 89.8% 90.9 89.8% 0.3% 10.0% 89.9
SBP 79.0% 0.6% 20.4% 79.3 90.6% 0.4% 9.0% 90.8 81.4% 0.2% 10.4% 89.5
AnP 77.0% 1.2% 21.8% 77.6 80.5% 0.6% 18.8% 90.9 81.4% 1.1% 17.5% 81.9
Table 2: Quantity of questions described in Section 4.2. The value $v_{ii^{\prime}}$ in the $i$th row and $i^{\prime}$th column represents the quantity of data that satisfies Theorem 4. Results prove that CoT has greater potential to significantly increase performance as scaling.
Qwen2.5-7B-Instruct
@@
DiP $\uparrow$ - 447 414 457 393 1711
CoT $\uparrow$ 423 - 374 416 361 1574
L2M $\uparrow$ 505 510 - 494 403 1912
SBP $\uparrow$ 599 601 564 - 429 2193
AnP $\uparrow$ 800 817 799 776 - 3192
Sum $\uparrow$ 2327 2375 2151 2143 1586 -
LLaMA-3-8B-Instruct
[…]
DiP $\uparrow$ - 816 794 459 513 2582
CoT $\uparrow$ 620 - 646 382 408 2056
L2M $\uparrow$ 639 677 - 432 393 2141
SBP $\uparrow$ 1316 1433 1398 - 923 5070
AnP $\uparrow$ 1243 1381 1380 871 - 4875
Sum $\uparrow$ 3818 4307 4218 2144 2237 -
@@ 4.1 CoT Has More Easy Questions and Fewer Hard Questions
We identify two primary reasons why CoT sometimes performs worse with lower sample sizes ($N$) but achieves better performance among these prompting approaches with larger $N$. The first reason relates to the distribution of question difficulty for CoT. CoT has more easy questions and fewer hard questions. When sampling with lower $N$, $\mathbf{P}_{i}$ still has a small probability of obtaining the right answer for hard questions, while the probability diminishes to zero as increasing $N$. This is the opposite of easy questions. Figure 39 shows an example of the accuracy changes on easy/hard questions. The prompting strategy with fewer hard questions and more easy questions will improve performance more rapidly when scaling. According to Theorem 1 to 3, we can calculate the extreme performance of $\mathbf{P}_{i}$ according to the difficulty proportion of questions, i.e., $\sum\limits_{x\in\mathfrak{D}}\lim\limits_{N\to+\infty}\!\!\!\!\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$. Table 1 summarizes the difficulty proportion of the questions and extreme performance for each $\mathbf{P}_{i}$ on each model. It can be observed that CoT has more easy questions and fewer hard questions, and can reach the best extreme performance on all models, thus making CoT gradually dominate as increasing $N$ even if it has a lower pass@1 accuracy.
@@ 4.2 CoT is Less Likely to be Affected by Wrong Answers
To quantify the uniformity of the erroneous answer distributions, we measure the KL divergence between them and the uniform distribution, where lower values indicate closer alignment with greater uniformity. Table 3 shows the average KL divergence on each $\mathbf{P}_{i}$ and LLM across all benchmarks. We can see that CoT has the most uniform distribution in most cases. While L2M achieves the lowest KL divergence on GLM-4-9B-Chat, CoT exhibits a marginal difference of merely 0.0003, indicating near-identical uniformity.
Table 3: Average KL divergence between the erroneous answer distribution and uniform distribution of each $\mathbf{P}_{i}$ across all tested 6 benchmarks on each LLM.
Qwen LLaMA GLM Phi Gemini GPT
DiP 0.0774 0.0830 0.0655 0.0770 0.0679 0.0649
CoT 0.0668 0.0829 0.0575 0.0678 0.0674 0.0624
L2M 0.0708 0.0857 0.0572 0.0724 0.0692 0.0647
SBP 0.0794 0.0952 0.0603 0.0821 0.0757 0.0647
AnP 0.1357 0.1154 0.1137 0.1097 0.1990 0.1593
As for why CoT has a more uniform probability distribution, we speculate this stems from its simpler approach that avoids imposing specific constraints or guidance on reasoning patterns, compared with other specifically designed prompting approaches. By minimizing explicit guidance, it preserves the model’s natural exploration of the solution space. While DiP is also simple, it fails to adequately activate the model’s step-by-step reasoning capabilities and free exploration potential. In contrast, more complex prompting strategies introduce explicit guidance mechanisms that constrain reasoning patterns and narrow search directions. This results in probability mass concentrating around specific potential correct solutions at the expense of broader exploratory behavior, ultimately leading to less uniform answer distributions.
$\mathbf{P}^{*}_{N}$ Sampling Time $N$
1 3 5 10 20 30 40 50 60 100 1000
Oracle L2M CoT CoT CoT SBP SBP SBP SBP SBP SBP SBP
Predicted (Ours) L2M CoT CoT CoT SBP SBP SBP SBP SBP SBP SBP
Correctness
@@ 5 Predicting Scaling Performance and $\mathbf{P}^{*}_{N}$
In practice, evaluating the test-time scaling performance requires significantly intensive resource consumption, especially with very large sampling time $N$. For pretraining, it is feasible to predict the train-time scaling performance based on the scaling law (Kaplan et al., 2020) through a series of low-cost experiments, while maintaining the model architecture largely unchanged and minimizing the risks associated with large-scale training. Similarly, we can also use the sample results of $\mathbf{P}_{i}$ with fewer $N$ to approximately get the distribution $\{p_{i,1},p_{i,2},\dots,p_{i,m}\}$ to predict the test-time scaling performance with larger $N$. Directly, one can utilize the multinomial distribution probability calculation formula (Equation 13 and 14) to calculate $\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ with enumeration or leverage numerical simulation to estimate. However, their computational complexities are both $O(N)$, and the former needs to traverse all situations and is difficult to operate. Therefore, we propose a method with the computational complexity $O(1)$ to quickly predict the scaling performance of majority voting for arbitrary $\mathbf{P}_{i}$, which can serve as the test-time scaling law for majority voting. It can also select the best prompting strategy $\mathbf{P}^{*}_{N}$ according to the predicted performances of each $\mathbf{P}_{i}$.
Here we omit the prompting index $i$ and input question $x$, and assume ${a}_{1}$ is the correct answer in the following. According to Khinchin’s Law of Large Numbers and Lindeberg-Levy Central Limit Theorem, when $N$ is large enough, each occurrence number $\mathbf{x}_{j}$ can be approximated by a normal distribution. Specifically, for $\mathbf{x}_{1}$, we have
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\mathbb{E}[Z]=N(p_{1}-p_{max}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$\mathbb{V}[Z]=N(p_{1}(1-p_{1})+p_{max}(1-p_{max})),$
@@ 7 Related Work
CoT series carefully design exemplars or 0-shot prompts to unleash the potential of step-by-step solving (Wei et al., 2022; Kojima et al., 2022; Zhang et al., 2023; Fu et al., 2023). Zhou et al. (2023); Dua et al. (2022); Khot et al. (2023) break down the question into smaller, more manageable subproblems. Madaan et al. (2023); Kim et al. (2023) force LLMs to self-evaluate and correct. Du et al. (2024); Liang et al. (2024); Chan et al. (2024); Liu et al. (2025); Huang et al. (2025) utilize multi-agent debate to collaborate reasoning. Yasunaga et al. (2024); Yu et al. (2024) guide LLMs to draw experience from analogous problems. Zheng et al. (2024); Gao et al. (2025) promote LLMs on abstract reasoning.
@@
Several studies have shown that scaling test-time compute optimally can be more effective than scaling model parameters (Snell et al., 2025; Open AI, 2024b). Snell et al. (2025) investigates the most effective test-time scaling approach with the basic fixed standard prompting given auxiliary resources, e.g., available datasets for training, specifically trained verifiers, and fine-tuned models. They analyze two kinds of test-time scaling approaches: 1) searching against verifier reward models, and 2) sequential revisions with specifically trained models. In contrast, our aim is, completely relying on the LLM itself, which prompting strategy is the most effective with the basic majority voting scaling. We provide new perspectives and theories to 1) understand, 2) predict, and 3) improve the scaling performance of different prompting strategies with the most basic test-time scaling setting.
@@ 8 Conclusion
We comprehensively study the behavior of various prompting strategies when scaling majority voting. Our experiments on 6 LLMs $\times$ 8 prompting strategies $\times$ 6 benchmarks consistently show that CoT has the potential to perform best as scaling. Theoretical analysis reveals that it benefits from fewer hard questions, more easy questions, and less susceptibility to incorrect answers, enabling more rapid performance gains. Additionally, our proposed method for predicting scaling performance offers a practical tool to select the optimal prompting strategy under given sampling time budgets. What’s more, we introduce two effective methods to further improve scaling performance.
We also extend experiments on two more challenging reasoning benchmarks, GPQA (Rein et al., 2024) and AIME (Mathematical Association of America, 2024), further verifying the generality of our findings and methods. Our combined method can achieve a significant boost, elevating accuracy from just over 30% (majority@100) to 75.7%. Please refer to Appendix E for more details.
@@ Acknowledgment
Appendix
In Appendix, we present broader related work, proofs for our theorems, detailed results on each benchmark, more discussions on improving the scaling performance, results of extended experiments, implementation details and prompts. The content structure is outlined as follows:
  • Appendix A - Broader Related Work
  • Appendix B - Proofs
  • Appendix C - Detailed Results
  • Appendix D - More Discussions on Improving the Scaling Performance
    • Appendix D.1 - Adaptively Scaling Based on the Difficulty
    • Appendix D.2 - Dynamically Choosing the Optimal $\mathbf{P}_{i}$
    • Appendix D.3 - Combining Adaptively Scaling and Dynamically Choosing the Optimal $\mathbf{P}_{i}$
  • Appendix E - Extended Experiments
  • Appendix F - Implementation Details and Prompts
@@ Appendix A Broader Related Work
Aggarwal et al. (2023); Li et al. (2024); Chen et al. (2024a) improve the reasoning efficiency with majority voting by adjusting the sampling time. Damani et al. (2025); Zhang et al. (2024b) learn to dynamically allocate resources under limited sampling time budgets. Chen et al. (2024b); Yue et al. (2024); Šakota et al. (2024) leverage multiple models with different prices to reduce cost while maintaining performance. Yang et al. (2025); Chen et al. (2024d); Yu et al. (2025) reduce the length of the thinking process to alleviate the overthinking issue, to achieve efficient reasoning.
@@
Jin et al. (2024b) studies the impact of reasoning step length of CoT. Wang et al. (2023a) studies what makes CoT prompting effective, indicating that being relevant to the query and correctly ordering the reasoning steps are more important. Feng et al. (2024); Cui et al. (2024) analyze the mechanism of CoT from a theoretical perspective. Sprague et al. (2025) points out that CoT helps mainly on math and symbolic reasoning by sorting and analyzing a large number of experimental results. Chen et al. (2024c) proposes a framework to quantify the reasoning boundary of CoT. Yang et al. (2024b) provides an in-context learning analysis of CoT. Chen et al. (2024a) investigates and analyzes the performance changes with more LLM calls. Chen et al. (2024e) proves that the failure probability of test-time scaling decays to zero exponentially or by a power law.
@@ Appendix B Proofs
$\!\!\!\!\!\!p_{1}^{k}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{t-1}^{k_{t-1}}p_{t}^{k_{t}+1}p_{t+1}^{k_{t+1}}\cdots p_{m}^{k_{m}}$.
It can be seen that case 1 and case 3 are also present in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$, whereas case 2 does not. We begin by considering case 1 and case 2. The terms in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ corresponding to case 1 $p_{1}^{k+1}p_{2}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}$ are shown in Equation 19, and no term in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ involves case 2. The terms in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ involving case 1 are shown in Equation 20. The terms in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$ involving case 2 $p_{1}^{k}p_{2}^{k}\cdots p_{s-1}^{k}p_{s}^{k+1}p_{s+1}^{k}\cdots p_{l}^{k}p_{l+1}^{k_{l+1}}p_{l+2}^{k_{l+2}}\cdots p_{m}^{k_{m}}$ are shown in Equation 21. Based on the fact of Equation 22, we can establish the inequality Equation 23, i.e., the terms corresponding to case 1 and case 2 in $\text{Pr}({a}_{1}|\mathbf{P}_{i};N+1)$ are greater than or equal to those in $(\sum_{j=1}^{m}{p_{j}})\cdot\text{Pr}({a}_{1}|\mathbf{P}_{i};N)$.
@@
Appendix C Detailed Results
Here we display the scaling performance of different prompting strategies on each LLM and benchmark under given sampling time $N$ and cost $O$, as shown in Figures 8 to 15. We find that, aside from CoT, DiP also exhibits superior performance compared to other complex prompting strategies on certain models and datasets, e.g., GPT-4o-mini on MATH. This also comes from the two reasons, i.e., DiP has more hard questions and easy questions, and a flat probability distribution of wrong answers on the specific dataset. This phenomenon is particularly prominent on powerful LLMs such as Gemini-1.5-Flash on GSM8K and GSM-Hard, where DiP and CoT exhibit comparable performance. Almost 83% of results satisfy that CoT or DiP performs best as significantly scaling. Besides, this trend is also observed on other prompting strategies on few datasets and models. This encourages us to fully unleash the potential of simple prompting strategies, and indicates that the scaling performance does not only depend on the prompting strategies’ pass@1 accuracy.